# Whisper-Flamingo: Integrating Visual Features into Whisper for Audio-Visual Speech Recognition and Translation

### v1

@@ Abstract
Audio-Visual Speech Recognition (AVSR) uses lip-based video to improve performance in noise. Since videos are harder to obtain than audio, the video training data of AVSR models is usually limited to a few thousand hours. In contrast, speech models such as Whisper are trained with hundreds of thousands of hours of data, and thus learn a better speech-to-text decoder. The huge training data difference motivates us to adapt Whisper to handle video inputs. Inspired by Flamingo which injects visual features into language models, we propose Whisper-Flamingo which integrates visual features into the Whisper speech recognition and translation model with gated cross attention. Our audio-visual Whisper-Flamingo outperforms audio-only Whisper on English speech recognition and En-X translation for 6 languages in noisy conditions. Moreover, Whisper-Flamingo is a versatile model and conducts all of these tasks using one set of parameters, while prior methods are trained separately on each language.
@@ 1 Introduction
Our experiments on the English (En) LRS3 video dataset [10] show that our novel audio-visual Whisper-Flamingo significantly outperforms the audio-only Whisper baseline in noise. Moreover, Whisper-Flamingo achieves competitive noise-robust results compared to prior audio-visual models. Next, we demonstrate Whisper’s multilingual capabilities by extending Whisper-Flamingo for En-X translation on the MuAViC dataset [11]. Our model performs En transcription and En-X translation into 6 other languages, while the previous audio-visual SOTA requires fine-tuning on each language separately. Once again, Whisper-Flamingo significantly outperforms audio-only Whisper in noise for both En transcription and En-X translation. Code and models at https://github.com/roudimit/whisper-flamingo
@@ 2 Method
The learnable parameters $\mathrm{\alpha_{xattn}}$ and $\mathrm{\alpha_{mlp}}$ are initialized to 0 so that the layers initially function as the identity since $\tanh(0)=0$. Through audio-visual fine-tuning, the model adjusts the weights of $\mathrm{\alpha_{xattn}}$ and $\mathrm{\alpha_{mlp}}$ and learns to attend to the visual features. We insert the gated cross attention layers in Whisper’s decoder in the beginning of each block, before the self-attention layer. We tried to insert them in other orders within the decoder blocks, but the performance was slightly worse. We also tried inserting them into the encoder, but the performance was significantly worse. Note that since the gated cross attention separately attends to the video features, the audio and video features can have different feature rates (for example, 50 Hz and 25 Hz).
@@
Table 3: Results for English transcription on LRS3. We report results on the original test set (Clean) and with babble noise added at 0-SNR (Noisy). A= Audio, AV= Audio-Visual. Noise dataset= dataset used to make babble noise. Hours of unlabeled / labeled data used to train each model are shown. Note that u-HuBERT [14] and AV-HuBERT [26] use a different noise file than us.
@@
Model Modalities Dataset A AV A AV Clean Noisy
@@
AV-BEST-RQ [6] AV NoiseX - 1759 - 433 2.1 6.8
AV2vec [15] AV MUSAN - 433 - 433 2.5 6.7
AV-HuBERT [26] AV LRS3† - 1759 - 433 1.4 5.8
u-HuBERT [14] AV LRS3† - 1759 - 433 1.3 4.6
@@
Adaptive AV [7] AV MUSAN - 400 680k 30 2.3 16.3
FAVA [6] AV NoiseX 1759 - - 433 1.7 6.6
FAVA-USM [6] AV NoiseX 12M - 5000 433 1.3 6.2
Our Audio-Only Whisper Baselines
Whisper-Large, Zero-shot (No Fine-Tuning) A LRS3 - - 680k - 2.1 20.8
Whisper-Large, Fine-tuned on LRS3 A LRS3 - - 680k - 2.3 11.7
Proposed Audio-Visual Fine-tuning Method
Whisper-Flamingo (Ours) AV LRS3 - 1759 680k 433 1.5 5.6
@@ 3 Experiments
Table 4: Results for English transcription on LRS3 and En-X Translation on MuAViC. Babble noise is added at 0-SNR (Noisy). One Model= the model translates to all languages with one set of parameters. Test Mod.= inference modalities (Text: T, audio: A, audio-visual: AV). Note that Bilingual AV-HuBERT [11] use a different noise file than us that was not publicly available.
@@ 3.1 Experimental Setup
To train our models, we use LRS3 [10] - the largest, publicly-available AVSR dataset in English (En), sourced from TED talks. We followed AV-HuBERT [4] to create a 433h training set, 1h validation set, and 1h test set. For En-X translation, we use the MuAViC [11] dataset which has translations of LRS3’s English text into 6 languages: Greek (El), Spanish (Es), French (Fr), Italian (It), Portuguese (Pt), and Russian (Ru).
@@
During training, we randomly add noise to the audio with a signal-to-noise ratio (SNR) of 0. Following prior work [26, 11], the “natural”, “music” and “babble” noise are sampled from the MUSAN dataset [42], and overlapping “speech” noise is sampled from LRS3 [10]. To select the best checkpoints, we monitor the highest token prediction accuracy on the noisy validation set every 1k steps. We report beam search decoding results with beam size 15. Following prior work [11], we use the Fairseq normalizer [43] to remove punctuation and lower-case text before calculating WER. For translation, we use SacreBLEU [44] with the default 13-a tokenizer to calculate BLEU [45].
@@ 3.3 Whisper-Flamingo English Speech Recognition
For our main experiments, we scale up to Whisper-Large. In the 3rd section of Table 3, we compare audio-only Whisper-Large zero-shot and fine-tuned on LRS3; the fine-tuned model outperforms the zero-shot model in noise (11.7% vs 20.8%). Fine-tuned Whisper-Large performs slightly worse than Whisper-Medium on clean audio (Table 2), but Whisper-Large performs better on noisy audio. The bottom part of Table 3 shows our audio-visual Whisper-Flamingo initialized from fine-tuned Whisper-Large. Compared to the audio-only baseline, Whisper-Flamingo significantly improves the noisy performance from 11.7% to 5.6% (52.1% relative improvement). It also improves the clean WER from 2.3% to 1.5%.
Table 3 also shows a comparison with prior audio-visual SSL methods and audio-visual fine-tuning methods on LRS3. Direct comparison in noisy conditions is challenging since different noise datasets were used to generate the babble noise. SSL methods AV-HuBERT [26] and u-HuBERT [14] used LRS3 to generate babble noise, but the noise file they generated was not publicly available. We followed their procedure to generate the noise, so our noisy conditions are similar but not identical. Compared with AV-HuBERT, Whisper-Flamingo achieves comparable clean performance (1.5% vs 1.4%) and slightly better noisy results (5.6% vs 5.8%), which shows that Whisper-Flamingo is effective at adapting Whisper to the visual features from AV-HuBERT. Moreover, a major advantage of Whisper-Flamingo over AV-HuBERT is improved translation performance (Section 3.4). The best noisy performance (4.6%) is reported by u-HuBERT; we would like to try it as a visual encoder for Whisper-Flamingo, but the weights are not publicly available. Finally, Whisper-Flamingo outperforms other methods in noise which adapt audio-only models through audio-visual fine-tuning [6, 7], including FAVA-USM [6] which was pre-trained on 12M hours of unlabeled audio [2]. However, the babble noise was generated from different datasets making results not strictly comparable.
@@ 5 Acknowledgments
We thank Alex H. Liu, Mohamed Anwar, and the reviewers for helpful discussion. This research was supported by the MIT-IBM Watson AI Lab and an NDSEG Fellowship to A.R.

### v2

@@ Abstract
Audio-Visual Speech Recognition (AVSR) uses lip-based video to improve performance in noise. Since videos are harder to obtain than audio, the video training data of AVSR models is usually limited to a few thousand hours. In contrast, speech models such as Whisper are trained with hundreds of thousands of hours of data, and thus learn a better speech-to-text decoder. The huge training data difference motivates us to adapt Whisper to handle video inputs. Inspired by Flamingo which injects visual features into language models, we propose Whisper-Flamingo which integrates visual features into the Whisper speech recognition and translation model with gated cross attention. Our models achieve state-of-the-art ASR WER (0.68%) and AVSR WER (0.76%) on LRS3. Audio-visual Whisper-Flamingo outperforms audio-only Whisper on English speech recognition and En-X translation for 6 languages in noisy conditions. Moreover, Whisper-Flamingo is versatile and conducts all of these tasks using one set of parameters, while prior methods are trained separately on each language.
@@ 1 Introduction
On the English (En) LRS3 video dataset [10], our models achieve State-of-the-Art (SOTA) ASR WER (0.68%) and AVSR WER (0.76%). Our novel audio-visual Whisper-Flamingo significantly outperforms the audio-only Whisper baseline in noise. Moreover, Whisper-Flamingo achieves competitive noise-robust results compared to prior audio-visual models. Next, we demonstrate Whisper’s multilingual capabilities by extending Whisper-Flamingo for En-X translation on the MuAViC dataset [11]. Our model performs En transcription and En-X translation into 6 other languages, while the previous audio-visual SOTA requires fine-tuning on each language separately. Once again, Whisper-Flamingo significantly outperforms audio-only Whisper in noise for both En transcription and En-X translation. Code and models at https://github.com/roudimit/whisper-flamingo
@@ 2 Method
The learnable parameters $\mathrm{\alpha_{xattn}}$ and $\mathrm{\alpha_{mlp}}$ are initialized to 0 so that the layers initially function as the identity since $\tanh(0)=0$. Through audio-visual fine-tuning, the model adjusts the weights of $\mathrm{\alpha_{xattn}}$ and $\mathrm{\alpha_{mlp}}$ and learns to attend to the visual features. We insert the gated cross attention layers in Whisper’s decoder in the beginning of each block, before the self-attention layer. We tried to insert them in other orders within the decoder blocks, but the performance was slightly worse. The full analysis is shown in the Appendix, Section 6.4. Note that since the gated cross attention separately attends to the video features, the audio and video features can have different feature rates (for example, 50 Hz and 25 Hz).
@@
Table 3: Results for English transcription on LRS3. We report results on the original test set (Clean) and with babble noise added at 0-SNR (Noisy). Hours of unlabeled & labeled audio-visual data used to train each model are shown. 433h=LRS3, 1,759h=LRS3+VoxCeleb2. Noise dataset= dataset used to make babble noise for testing. Noisy WER from methods using different noise datasets are not directly comparable. Note that u-HuBERT [14], AV-HuBERT [26], and CMA [27] use a different noise file than us. *=noisy results were not reported using 0-SNR.
@@
Model Dataset Unlabeled Labeled Clean Noisy Clean Noisy
Audio-Visual methods trained only on transcribed videos
VIT 3D [17] NoiseX - 90k 1.6 6.1 1.6 2.9
LP Conformer [28] NoiseX - 100k - - 0.9 1.9
AutoAVSR [21] NoiseX - 1,902 / 3,448 1.0 / 1.0 * 1.0 / 0.9 *
Fast Conformer [29] NoiseX - 435 / 3,116 1.6 / 0.7 * 1.7 / 0.8 *
@@
AV2vec [15] MUSAN 433 433 2.7 19.5 2.5 6.7
AV-BEST-RQ [6] NoiseX 1,759 433 - - 2.1 6.8
BRAVEn [30] NoiseX 2,649 433 1.1 - - -
AV-HuBERT [26] LRS3† 1,759 433 1.6 15.7 1.4 5.8
u-HuBERT [14] LRS3† 1,759 433 1.4 19.3 1.3 4.6
CMA [27] LRS3† 1,759 433 - - 1.5 4.4
@@
Adaptive AV [7] MUSAN 400 30 - - 2.3 16.3
FAVA-USM [6] NoiseX - 433 - - 1.3 6.2
Llama-AVSR [31] NoiseX 1,759 433 / 1,759 1.1 / 0.81 12.3 / - 0.95 / 0.77 4.2 / -
Our audio-only Whisper Zero-shot baseline
Whisper-Large, Zero-shot, Beam 1 LRS3 - - 2.3 23.3 - -
$\hookrightarrow$ w/ Beam search 15 LRS3 - - 2.1 20.8 - -
Our models trained without noise
Whisper-Large, Fine-tuned w/o noise, Beam 1 LRS3 - 433 / 1,759 1.0 / 0.68 21.6 / 23.6 - -
$\hookrightarrow$ w/ Beam search 15 LRS3 - 433 / 1,759 1.3 / 1.5 23.1 / 24.1 - -
Whisper-Flamingo, Fine-tuned w/o noise, Beam 1 LRS3 1,759 433 / 1,759 - - 1.1 / 0.76 13 / 14.9
$\hookrightarrow$ w/ Beam search 15 LRS3 1,759 433 / 1,759 - - 1.1 / 1.8 12.8 / 15.9
Our models trained with noise
Whisper-Large, Fine-tuned w/ noise, Beam 1 LRS3 - 433 / 1,759 1.1 / 0.85 12.4 / 11.9 - -
$\hookrightarrow$ w/ Beam search 15 LRS3 - 433 / 1,759 2.3 / 2.0 11.7 / 11.1 - -
Whisper-Flamingo, Fine-tuned w/ noise, Beam 1 LRS3 1,759 433 / 1,759 - - 1.0 / 0.86 7.2 / 6.1
$\hookrightarrow$ w/ Beam search 15 LRS3 1,759 433 / 1,759 - - 1.5 / 2.0 5.6 / 5.6
@@ 3 Experiments
Table 4: Results for English transcription on LRS3 and En-X Translation on MuAViC. Babble noise is added at 0-SNR (Noisy). One Model= the model translates to all languages with one set of parameters. Test Mod.= inference modalities (Text: T, audio: A, video: V audio-visual: AV). Note that Bilingual AV-HuBERT [11] use a different noise file than us that was not publicly available so the results in noisy conditions are not directly comparable.
@@
Speech-to-Text Translation (Video only, no audio)
VSP-LLM [39] V - - - 22.7 22.3 17.9 18.7 - -
@@ 3.1 Experimental Setup
To train our models, we use LRS3 [10] - the largest, publicly-available AVSR dataset in English (En), sourced from TED talks. We followed AV-HuBERT [4] to create a 433h training set, 1h validation set, and 1h test set. We also combined the LRS3 training videos with 1,326h of English videos from VoxCeleb2 [40] for training. Transcripts for the VoxCeleb2 videos were obtained from Whisper Large-v2 [41]. For En-X translation, we use the MuAViC [11] dataset which has translations of LRS3’s English text into 6 languages: Greek (El), Spanish (Es), French (Fr), Italian (It), Portuguese (Pt), and Russian (Ru).
@@
We train our models in two conditions: without noise (clean) and with noise (noisy). For clean training, we do not add any noise to the audio. For noisy training, we randomly add noise to the audio with a signal-to-noise ratio (SNR) of 0. Following prior work [26, 11], the “natural”, “music” and “babble” noise are sampled from the MUSAN dataset [49], and overlapping “speech” noise is sampled from LRS3 [10]. To select the best checkpoints, we monitor the highest token prediction accuracy on the clean or noisy validation set every 1k steps. Following prior work [11], we use the Fairseq normalizer [50] to remove punctuation and lower-case text before calculating WER. For translation, we use SacreBLEU [51] with the default 13-a tokenizer to calculate BLEU [52].
@@ 3.3 Whisper-Flamingo English Speech Recognition
For our main experiments in Table 3, we use Whisper-Large. We report results using beam search with beam size 1 and 15. In noisy conditions, we use babble noise at 0-SNR constructed following AV-HuBERT [26] by adding audio from 30 speakers from LRS3. Results for additional noise types and SNR levels are shown in the Appendix, Section 6.2.
Training on clean audio (without noise). Compared to zero-shot audio-only Whisper-Large, fine-tuning without noise improves the clean ASR WER from 2.1% to 1.0% (using LRS3 433h) and 0.68% (using LRS3+VoxCeleb2 1,759h). Our fine-tuned Whisper-Large achieves SOTA ASR on LRS3 (0.68%) and matches the previous SOTA of 0.7% from Fast Conformer [29]. We then use our fine-tuned Whisper models to initialize audio-visual Whisper-Flamingo and achieve 1.1% / 0.76% AVSR WER using LRS3 433h / LRS3+VoxCeleb2 1,759h respectively. Our audio-visual Whisper-Flamingo achieves SOTA AVSR on LRS3 (0.76%) and matches the current SOTA from Llama-AVSR (0.77%) using LRS3+VoxCeleb2 and Fast Conformer (0.8%) using additional data. Notably, our method only uses 2.5B parameters, while Llama-AVSR uses over 8B parameters. Also, our method achieves better AVSR performance than LP Conformer (0.9%) which was trained on 100k videos. Fine-tuning Whisper on clean audio does not improve the noisy WER, while training Whisper-Flamingo on clean audio improves the noisy WER.
Training on noisy audio. Compared to zero-shot audio-only Whisper-Large, fine-tuning with noise improves the noisy ASR WER from 20.8% to 11.7% (using LRS3 433h) and 11.1% (using LRS3+VoxCeleb2 1,759h). The clean ASR WER is slightly worse compared to Whisper fine-tuned without noise, however, the results are close (0.85% vs 0.68%). We then use our fine-tuned Whisper models to initialize audio-visual Whisper-Flamingo and significantly improve the noisy WER to 5.6% / 5.6% AVSR WER using LRS3 433h / LRS3+VoxCeleb2 1,759h respectively (49.5% relative noisy WER improvement compared to fine-tuned audio-only Whisper). The clean AVSR WER is slightly worse compared to Whisper-Flamingo trained without noise, however, the results are close (0.86% vs 0.76%).
Comparing noisy results to SOTA. Table 3 also shows a comparison with prior audio-visual SSL methods and audio-visual fine-tuning methods on LRS3. Direct comparison in noisy conditions is challenging since different noise datasets were used to generate the babble noise. SSL methods AV-HuBERT [26], u-HuBERT [14], and CMA [27] used LRS3 to generate babble noise, but the noise file they generated was not publicly available. We followed their procedure to generate the noise, so our noisy conditions are similar but not identical. Compared with AV-HuBERT, Whisper-Flamingo achieves better clean performance (0.86% vs 1.4%) and slightly better noisy results (5.6% vs 5.8%), which shows that Whisper-Flamingo is effective at adapting Whisper to the visual features from AV-HuBERT. Moreover, a major advantage of Whisper-Flamingo over AV-HuBERT is improved translation performance (Section 3.4). The best noisy performances are reported by u-HuBERT and CMA; we would like to try them as visual encoders for Whisper-Flamingo, but the weights are not publicly available. Finally, Whisper-Flamingo outperforms other methods in noise which adapt audio-only models through audio-visual fine-tuning [6, 7], including FAVA-USM [6] which was pre-trained on 12M hours of unlabeled audio [2]. However, the babble noise was generated from different datasets making results not strictly comparable.
@@ 3.4 Whisper-Flamingo En-X Speech Translation
For these experiments, we fine-tune Whisper and train Whisper-Flamingo with noise. We report results with a beam size of 15.
@@
Comparison with VSP-LLM. Finally, we compare with VSP-LLM [39], a recent LLM-based approach which uses features from AV-HuBERT as input to an LLM to perform lip-reading and translation using video inputs only (no audio). Even with noise in the audio, our Whisper-Flamingo (Large) outperforms VSP-LLM on all 4 languages evaluated, which shows the advantage of using both audio and video as inputs instead of just video. Also, our model only has 2.5B parameters compared with VSP-LLM’s 7B parameters.
@@ 5 Acknowledgments
We thank Alex H. Liu, Mohamed Anwar, and the reviewers for helpful discussion. We thank Videet Mehta for help with the experiments in Table A2. This research was supported by the MIT-IBM Watson AI Lab and an NDSEG Fellowship to A.R.